A Bibliography of Publications of W. J. Cody

Nelson H. F. Beebe Department of Mathematics

University of Utah Salt Lake City, UT 84112

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,  beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

12 August 2025 Version 1.57

Abstract

This bibliography records publications of W. J. Cody.

Title word cross-reference

−1/2 [11, 15]. 1/2 [11, 15]. 3/2 [11, 15]. [0, ∞) [22]. E [4, 6, 3, 3]. e^{−x} [22]. E_{1}(x) [18]. Ei(x) [21]. K [4, 6, 3, 3].

0.5e [60].

10th [127]. 11/780 [74]. 1st [118].

20th [119].

360 [8, 10]. 3600 [1].

82-63 [60].

Accuracy [87, 96, 25, 38, 64]. ACM [112]. Address [94]. Algorithm [62, 79, 84, 89, 97, 98, 99, 28]. algorithms [12]. Alliant [77]. Alternative [74]. Analysis [54, 88, 115, 47]. Annual [112, 118]. Applications [114, 22]. approach [40]. Approximation [29, 43, 100]. Approximations [4, 6, 9, 14, 21, 26, 101, 7, 11, 15, 16, 17, 18, 19, 22, 24, 27, 31, 35]. April [108, 116, 114]. Architectural [119]. Architecture [118]. Ardent [85]. Argonne [124]. Arithmetic [37, 58, 66, 72, 84, 94, 98, 127, 126, 53, 60, 63]. Art [124]. Aspect [13]. AST [90]. Atlantic [108]. atoms [2]. August [112, 118, 122]. available [81].

Balance [76]. Basic [56]. basics [48, 51]. Bessel [46, 50, 62, 86]. Between [122]. Binary [84, 98]. Boston [112, 109]. Boulder [121, 122]. Brook [117]. BSD [74]. Bureau [121].

CA [119]. California [119]. Carolina [113]. CDC [1]. CDC-3600 [1]. CELEFUNT [93, 97]. Certain [86]. certify [34]. challenge [45]. Chapel [113]. Characteristics [32, 37]. Chebyshev [4, 5, 6, 7, 9, 11, 15, 16, 17, 18, 19, 21, 22, 24, 26, 27, 31, 35]. City [108]. Cody [106, 105, 107]. collaborative [34]. Collins [118]. Colorado [118, 121, 122]. COMPCON [119]. Complementary [83, 92]. Complete [4, 5, 6, 3]. Complex [93, 97]. Computation [32, 46, 122]. Computational [124, 48, 51, 56]. Computer [108, 109, 110, 116, 123, 113, 127, 102, 128, 126]. Computing [14, 125, 104]. concepts [56]. conduction [22]. Conference [112, 108, 109, 110, 114, 121, 122, 115]. Construction [39]. Contained [30, 20]. Corrigenda [6]. Corrigendum [11, 17]. Coulomb [26, 28]. Critique [10].

Dawson [27]. design [12]. Desirable [32]. Determine [79]. Development [103, 42]. Digest [119]. Dirac [11, 15]. disseminate [34]. Document [60]. Double [1]. Double-Precision [1]. Draft [60]. Drivers [89, 99]. Dundee [115]. Dynamic [37]. Dynamically [79].

Economical [13]. Editor [13]. Effort [65, 34, 70, 71]. elastic [2]. ELEFUNT [75, 76, 77, 85, 90]. Elementary [33, 52, 67, 93, 97, 61]. Elliptic [4, 5, 6, 3]. Encore [75]. Encyclopedia [102, 128]. Environments [104]. equations [19]. Error [83, 92, 24]. Evaluating [101, 121]. Evaluation [82, 83, 86, 92, 95, 101, 36]. Expansions [5]. Exponential [14, 21, 18]. exponentiation [20].

Fall [110]. February [119]. Fermi [11, 15]. Fifteenth [123]. First [62, 50]. Floating [37, 49, 54, 57, 58, 66, 72, 80, 84, 88, 98, 25, 38, 53, 60, 63, 64]. Floating-Point [37, 54, 58, 66, 80, 84, 88, 98, 57, 72, 53, 60, 63]. Fort [118]. Fortran [90, 8, 10, 99, 76, 77, 85]. Fractional [46]. France [127]. Francisco [119]. Free [60]. Fresnel [16]. Function [9, 14, 41, 55, 59, 82, 87, 95, 96, 99, 101, 23, 24, 31, 35, 61, 68, 78]. Functions [28, 29, 33, 46, 52, 62, 67, 83, 84, 86, 89, 92, 93, 97, 98, 42, 50, 69, 73]. FUNPACK [41, 55, 68]. FX [77]. FX/8 [77]. FX/Fortran [77].

Gamma [9, 82, 95]. Generalization

[58, 63]. Grenoble [127].

H [8, 10, 124]. Hardware [32]. Harvard [116]. heat [22]. Held [121, 124, 114, 113, 120]. Hill [113]. honor [124]. Horizons [119]. Hotel [119]. Houston [123]. hydrogen [2].

IBM [8, 10]. IEEE

[127, 49, 58, 60, 63, 84, 98]. IFIP [122].

Illinois [117]. Impact [49].

Implementation [59]. independent

[60, 66, 72]. influence [12]. Informal [124]. Institute [116, 114]. Integral [21, 91, 18, 27]. Integrals [4, 5, 6, 3, 11, 15, 16]. Interface [116, 123]. International [120]. interpolation [7, 17]. Interval [125]. interview [105]. Italy [120]. IV [8, 10].

J [105]. Jack [119]. James [124]. January [121]. Jim [107]. Joint [108, 109, 110]. July [115]. June [117, 113, 127].

Keynote [94]. Kind [46, 62, 50].

Laboratory [124]. Languages [122]. Large [104]. Large-Scale [104]. length [66, 72]. Letter [13]. Libraries [39, 107]. Library [8, 10, 74]. linear [19]. Logarithm [9]. Long [94]. Loughborough [114].

MA [109]. MACHAR [79, 81]. Machine [47, 79, 12]. Manual [52]. March [123]. Massachusetts [116]. Mathematical [65, 103, 120, 121, 111, 34, 36, 44, 70, 71, 107]. Mathematics [124, 114, 48, 51]. May [109]. Methodologies [120]. Methods [101, 125, 113]. Microcomputers [118, 104]. minicomputers [45]. models [57]. Modified [62, 50]. Mountain [118]. Multimax [75].

National [124, 121]. NATS [34, 40]. Natural [9]. Nevada [110]. Ninth [116]. NJ [108]. Normal [91]. North [113]. Note [14]. November [110, 120]. NS32000 [76]. Number [25, 38, 64]. Numerical [37, 39, 49, 117, 114, 122, 115, 12, 45, 47].

Oak [117]. Obituary [106]. Observations [65]. Order [46]. orders [11, 15]. overview [42].

P854 [60]. P854/82 [60]. P854/82-63 [60]. Package [41, 55, 89, 93, 97, 99, 68, 78]. Papers [119]. Parameters [79, 47, 57]. Performance [8, 23, 82, 83, 86, 92, 95]. perspective [70, 71]. Phase [26]. Pioneer [107]. Point [37, 49, 54, 58, 66, 80, 84, 88, 98, 25, 38, 53, 57, 60, 63, 72, 64]. Polynomial [5, 29, 101]. Polynomials [13]. Portability [117]. Portable [89, 93, 97, 99, 78]. positrons [2]. Power [30]. Practical [29]. Practice [80]. Precision [1]. Preliminary [50]. Problems [120, 22]. procedures [61]. Proceedings [112, 116, 114, 123, 113, 118, 121, 122, 115, 124, 127]. Production [120]. Program [113]. Programming [121, 122]. Programs [82, 83, 86, 87, 92, 95, 96]. Proposals [54, 88]. Proposed [49, 58, 66, 72, 60, 63]. psi [35].

quality [40].

Radix [66, 72, 60]. Radix- [66, 72].

Radix-Free [60]. radix-independent [60]. Rational [7, 15, 18, 19, 24, 29, 22, 11, 17]. Real [46, 82, 95]. Regular [28]. Related [82, 95]. Relationship [122]. Reliability [125]. Remark [28]. report [50]. Reprint [64]. Results [75, 76, 77, 85, 90]. Riemann [31]. Road [94]. Robustness [44]. Rocky [118]. Role [125]. Root [1]. Routines [30, 99, 101, 68].

S22 [28]. San [119]. Scale [104]. scattering [2]. Science [116, 123, 102, 128]. Scientific [32, 125]. Second [46, 70, 71]. Self [20, 30]. Self-Contained [30, 20]. Seminar [120]. sensible [53]. September [124, 118]. Sequence [62]. Sequent [76, 90]. Series [87, 96]. Shift [26]. slow [2]. Software [33, 48, 49, 51, 52, 59, 65, 67, 69, 73, 117, 103, 114, 118, 120, 111, 34, 36, 42, 44, 45, 50, 56, 61, 70, 71, 107, 40]. Sorrento [120]. Sources [103]. SPECFUN [78, 89, 99]. Special [41, 55, 89, 99, 42, 68, 69, 73, 78]. Spring [108, 109, 119]. Square [1]. Standard [49, 54, 58, 66, 72, 84, 88, 98, 60, 63]. Standards [80, 94, 121, 57]. State [124]. Static [37]. Statistical [25, 38, 64]. Statistics [116, 8, 123]. Study [25, 38, 64]. Subcommittee [60]. Subroutine [39, 79]. Subroutines [41, 55, 23]. Support [84, 98]. Survey [29]. Symmetry [90]. Symposium [116, 124, 123, 113, 118, 127]. System [8, 10]. System/360 [8, 10]. Systems [118, 25, 38, 64].

Tables [3]. Tar [119]. Taylor [87, 96]. TC2 [122]. Techniques [121]. Technology [116, 114]. Test [75, 76, 77, 85, 87, 89, 90, 93, 96, 97, 99, 61, 113]. Testing [59, 23]. Texas [123]. Theory [43, 80, 100]. thoughts [70, 71]. Titan [85]. Transportable [61].

University [116, 114, 113]. UNIX [74, 85]. USA [122]. Use [87, 96]. Using [85, 7, 17, 19].

V1.8.0 [90]. V2.5.3 [76]. VAX [74]. Vegas [110]. Version [77, 81]. VLSI [119].

W. [105]. Wave [28]. Wilkinson [124]. William [106]. Word [66, 72]. Word-length-independent [66, 72]. Working [60, 122]. Workshop [117].

X1.4 [75]. XXX [84, 89].

zeta [31].

References

Cody:1964:DPS

[1] W. J. Cody, Jr. Double-precision square root for the CDC-3600. Communications of the Association for Computing Machinery, 7(12):715–718, December 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cody:1964:ESS

[2] W. J. Cody, Joan Lawson, Sir Harrie Massey, F.R.S., and K. Smith. The elastic scattering of slow positrons by hydrogen atoms. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 278:479–489, 1964. CODEN PRLAAZ. ISSN 0080-4630.

Fisherkeller:1964:TCE

[3] M. A. Fisherkeller and W. J. Cody, Jr. Tables of the complete elliptic integrals K, K′, E, and E′. Technical Memo ANL AMD 71, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, 1964. 14 pp. See review by John W. Wrench in Mathematics of Computation, 19(89–92), 342, 1965.

Cody:1965:CAC

[4] W. J. Cody, Jr. Chebyshev approximations for the complete elliptic integrals K and E. Mathematics of Computation, 19(89–92):105–112, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://

www.jstor.org/stable/2004103. See

also [6].

Cody:1965:CPE

[5] W. J. Cody, Jr. Chebyshev polynomial expansions of complete elliptic integrals. Mathematics of Computation, 19 (89–92):249–259, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://

www.jstor.org/stable/2003350.

Cody:1966:CCA

[6] W. J. Cody, Jr. Corrigenda: “Chebyshev approximations for the complete elliptic integrals K and E”. Mathematics of Computation, 20(93):207, January 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See [4].

Cody:1966:RCA

[7] W. J. Cody and J. Stoer. Rational Chebyshev approximations using interpolation. Numerische Mathematik, 9:177–188, 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See also [17].

Clark:1967:PSF

[8] N. A. Clark, W. J. Cody, K. E. Hillstrom, and E. A. Thieleker. Performance statistics of the FORTRAN IV (H) library for the IBM System/360. Technical Report ANL-7231, Argonne National Laboratory, Argonne, IL, USA, 1967. Reprinted in SHARE Secretary Distribution, SDD 169, C4473, pp. 12 46.

Cody:1967:CAN

[9] W. J. Cody, Jr. and K. E. Hillstrom. Chebyshev approximations for the natural logarithm of the gamma function. Mathematics of Computation, 21(98): 198–203, April 1967. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842

(electronic). URL http://www.jstor.
org/stable/2004160.

Cody:1967:CFI

[10] W. J. Cody. Critique of the FORTRAN IV (H) library for the IBM System/360. SHARE Secretary Distribution, SSD 169 (C4473):4–11, 1967.

Cody:1967:CRC

[11] W. J. Cody and Henry C. Thacher, Jr. Corrigendum: “Rational Chebyshev approximations for Fermi–Dirac integrals of orders −1/2, 1/2, and 3/2”. Mathematics of Computation, 21(99): 525, July 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.jstor.org/
stable/2003289. See [15].

Cody:1967:IMD

[12] W. J. Cody. The influence of machine design on numerical algorithms. In AFIPS SJCC ’67 [108], pages 305–309. LCCN TK7885.A1 J6 1967.

Cody:1967:LEA

[13] William J. Cody, Jr. Letter to the Editor: Another aspect of economical polynomials. Communications of the Association for Computing Machinery, 10(9):531, September 1967. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [101].

Cody:1967:NCA

[14] W. J. Cody and Anthony Ralston. A note on computing approximations to the exponential function. Communications of the Association for Computing Machinery, 10(1):53–55, January 1967. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cody:1967:RCA

[15] W. J. Cody and Henry C. Thacher, Jr. Rational Chebyshev approximations for Fermi–Dirac integrals of orders −1/2, 1/2, and 3/2. Mathematics of Computation, 21(97):30–40, January 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2003468. See also [11].

Cody:1968:CAF

[16] W. J. Cody. Chebyshev approximations for the Fresnel integrals. Mathematics of Computation, 22(102):450– 453 (plus microfiche supplement), April 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cody:1968:CRC

[17] W. J. Cody and J. Stoer. Corrigendum: “Rational Chebyshev approximations using interpolation”. Numerische Mathematik, 12:230, 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See [7].

Cody:1968:RCAa

[18] W. J. Cody and H. C. Thacher, Jr. Rational Chebyshev approximations for the exponential integral E_{1}(x). Mathematics of Computation, 22(103):641–649, July 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cody:1968:RCAb

[19] W. J. Cody, W. Fraser, and J. F. Hart. Rational Chebyshev approximations using linear equations. Numerische Mathematik, 12:242–251, 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Clark:1969:SCE

[20] N. W. Clark and W. J. Cody. Selfcontained exponentiation. In AFIPS FJCC ’69 [110], pages 701–706. LCCN TK7885.A1 J6 1969.

Cody:1969:CAE

[21] W. J. Cody and Henry C. Thacher, Jr. Chebyshev approximations for the exponential integral Ei(x). Mathematics of Computation, 23(106):289–303, April 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cody:1969:CRA

[22] W. J. Cody, G. Meinardus, and R. S. Varga. Chebyshev rational approximations to e^{−x} on [0, ∞) and applications to heat conduction problems. J. Approx. Theory, 2(??):50–65, ?? 1969. CODEN JAXTAZ. ISSN 0021-9045.

Cody:1969:PTF

[23] W. J. Cody. Performance testing of function subroutines. In AFIPS SJCC ’69 [109], pages 759–763. LCCN TK7885.A1 J6 1969.

Cody:1969:RCA

[24] W. J. Cody, Jr. Rational Chebyshev approximations for the error function.

Mathematics of Computation, 23(107): 631–637, July 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.jstor.org/
stable/2004390.

Kuki:1969:SSA

[25] H. Kuki and W. J. Cody. A statistical study of the accuracy of floating point number systems. Communications of the Association for Computing Machinery, 16(4):223–230, April 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cody:1970:CAC

[26] W. J. Cody and K. E. Hillstrom. Chebyshev approximations for the Coulomb phase shift. Mathematics of Computation, 24(111):671–677, July 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cody:1970:CAD

[27] W. J. Cody, Kathleen A. Paciorek, and Henry C. Thacher, Jr. Chebyshev approximations for Dawson’s integral. Mathematics of Computation, 24(109): 171–178, January 1970. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cody:1970:RAC

[28] W. J. Cody and Kathleen A. Paciorek. Remark on Algorithm 292 [S22]: Regular Coulomb wave functions. Communications of the Association for Computing Machinery, 13(9):573, September 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cody:1970:SPR

[29] W. J. Cody. A survey of practical rational and polynomial approximation of functions. SIAM Review, 12(3):400– 423, July 1970. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). Reprinted in Studies in Applied Mathematics, 6, SIAM, 1970, pp. 86– 109.

Clark:1971:SCP

[30] N. W. Clark, W. J. Cody, and H. Kuki. Self-contained power routines. In Mathematical Software [111], pages 399–415. ISBN 0-12-587250-X. LCCN QA1 .M26.

Cody:1971:CAR

[31] W. J. Cody, K. E. Hillstrom, and Henry C. Thatcher, Jr. Chebyshev approximations for the Riemann zeta function. Mathematics of Computation, 25(115):537–547, July 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cody:1971:DHC

[32] W. J. Cody. Desirable hardware characteristics for scientific computation. ACM SIGNUM Newsletter, 6(1):16–31, January 1971. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Cody:1971:SEF

[33] W. J. Cody. Software for the elementary functions. In Mathematical Software [111], pages 171–186. ISBN 0-12-587250-X. LCCN QA1 .M26.

Boyle:1972:NCE

[34] J. M. Boyle, W. J. Cody, W. R. Cowell, B. S. Garbow, Y. Ikebe, C. B. Moler, and B. T. Smith. NATS, a collaborative effort to certify and disseminate mathematical software. In ACM ’72 [112], pages 630–635. LCCN TK 7885 A84p 1972. Two volumes.

Cody:1973:CAP

[35] W. J. Cody, Jr., Anthony J. Strecok, and Henry C. Thacher, Jr. Chebyshev approximations for the psi function. Mathematics of Computation, 27(21): 123–127, January 1973. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842

(electronic). URL http://www.jstor.
org/stable/2005253.

Cody:1973:EMS

[36] W. J. Cody. The evaluation of mathematical software. In Hetzel [113], pages 121–133. ISBN 0-13-729624-X. LCCN QA76.P69.

Cody:1973:SDN

[37] William J. Cody, Jr. Static and dynamic numerical characteristics of floating-point arithmetic. IEEE Transactions on Computers, C-22(6):598– 601, June 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=5009112;   http:
//www.acsel-lab.com/arithmetic/
arith2/papers/ARITH2_Cody.pdf.

Kuki:1973:SSA

[38] H. Kuki and W. J. Cody. A statistical study of the accuracy of floating point number systems. Communications of the Association for Computing Machinery, 16(4):223–230, April 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cody:1974:CNS

[39] W. J. Cody. The construction of numerical subroutine libraries. SIAM Review, 16(1):36–46, January 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Smith:1974:NAQ

[40] B. T. Smith, J. M. Boyle, and W. J. Cody. The NATS approach to quality software. In Evans [114], pages 393– 405. ISBN 0-12-243750-0. LCCN QA297 .S591.

Cody:1975:FPS

[41] W. J. Cody. The FUNPACK package of special function subroutines. ACM Transactions on Mathematical Software, 1(1):13–25, March 1975. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Cody:1975:OSD

[42] W. J. Cody. An overview of software development for special functions. In Watson [115], pages 38–48. ISBN 3-540-07610-7 (print), 3-540-38129-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA297; QA3 .L28 no. 506.

Cody:1976:AT

[43] W. J. Cody. Approximation theory. In Ralston and Meek [102], pages 87–90. ISBN 0-88405-321-0. LCCN QA76.15 .E56 1976. US$60.00.

Cody:1976:RMS

[44] W. J. Cody. Robustness in mathematical software. In Hoaglin and Welsch [116], pages 76–78. ISBN 0-87150-237-2. LCCN QA 276 A1 I53 1976.

Cody:1977:CNS

[45] W. J. Cody. The challenge in numerical software for minicomputers. In IEEE RMSS ’77 [118], pages 1–23. LCCN QA76.5 .R535 1977a.

Cody:1977:CRF

[46] W. J. Cody, Rose M. Motley, and L. Wayne Fullerton. The computation of real fractional order Bessel functions of the second kind. ACM Transactions on Mathematical Software, 3(3): 232–239, September 1977. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Cody:1977:MPN

[47] W. J. Cody. Machine parameters for numerical analysis. In Cowell [117], pages 49–67. ISBN 0-387-08446-0. LCCN QA297 .W65 1976.

Cody:1978:SBC

[48] W. J. Cody. Software basics for computational mathematics. Technical Memo ANL AMD 328, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, November 1978. Draft form of [51].

Cody:1979:IPI

[49] W. Cody. Impact of the proposed IEEE floating point standard on numerical software. ACM SIGNUM Newsletter, 14 (special issue):29–30, October 1979. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Cody:1980:PRS

[50] W. J. Cody. Preliminary report on software for modified Bessel functions of the first kind. Amd tech. memo tm-357, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, 1980. ?? pp.

Cody:1980:SBC

[51] W. J. Cody. Software basics for computational mathematics. ACM SIGNUM Newsletter, 15(2):18–29, June 1980. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic). Reprinted from “Program Directions for Computational Mathematics”, R. Huddleston (ed.), DOE, June 1979.

Cody:1980:SME

[52] William J. Cody, Jr. and William Waite. Software Manual for the Elementary Functions. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1980. ISBN 0-13-822064-6. x + 269 pp. LCCN QA331 .C635 1980.

Cody:1980:TSF

[53] W. J. Cody. Towards sensible floatingpoint arithmetic. In IEEE COMPCON Spring ’80 [119], pages 488–490. LCCN TK7885.A1 C53 1980.

Cody:1981:APF

[54] William J. Cody, Jr. Analysis of proposals for the floating-point standard. Computer, 14(3):63–68, March 1981. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cody:1981:FPS

[55] W. J. Cody. FUNPACK—A package of special function subroutines. Technical Report TM-385, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, April 1981. ii + 14 pp. Republished in [68].

Cody:1982:BCC

[56] W. J. Cody. Basic concepts for computational software. In Messina and Murli [120], pages 1–23. ISBN 0-387-11603-6 (New York), 3-540-11603-6 (Berlin). LCCN QA76.95 .P76 1982.

Cody:1982:FPM

[57] W. J. Cody. Floating-point parameters, models, and standards. In Reid [122], pages 51–69. ISBN 0-444-86377-X. LCCN QA297 .I34 1981.

Cody:1982:GPI

[58] W. J. Cody. A generalization of the proposed IEEE standard for floating-point arithmetic. Technical Report ??, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, 1982. 20 pp. Republished in [63].

Cody:1982:ITF

[59] W. J. Cody. Implementation and testing of function software. In Messina and Murli [120], pages 24–47. ISBN 0-387-11603-6 (New York), 3-540-11603-6 (Berlin). LCCN QA76.95 .P76 1982.

Cody:1982:PRI

[60] W. J. Cody, Chairman. A proposed radix-independent standard for floatingpoint arithmetic, draft 0.5e, IEEE radixfree floating point subcommittee working document P854/82-63. Technical report, IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1982. ?? pp.

Cody:1982:TTP

[61] W. J. Cody. Transportable test procedures for elementary function software. In Mulvey [121], pages 236–247. ISBN 0-387-11495-5. LCCN QA402.5 .E94 1982.

Cody:1983:ASM

[62] W. J. Cody. Algorithm 597: Sequence of modified Bessel functions of the first kind. ACM Transactions on Mathematical Software, 9(2):242–245, June 1983. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Cody:1983:GPI

[63] W. J. Cody. A generalization of the proposed IEEE standard for floatingpoint arithmetic. In Gentle [123], pages 133–139. ISBN 0-444-86688-4. LCCN QA276.4 .S95 1983. Republication of [58].

Kuki:1983:SSA

[64] H. Kuki and W. J. Cody. A statistical study of the accuracy of floating point number systems (reprint). Communications of the Association for Computing Machinery, 26(1):79–83, January 1983. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Reprint of [38].

Cody:1984:OMS

[65] W. J. Cody. Observations on the mathematical software effort. In Cowell [103], pages 1–19. ISBN 0-13-823501-5. LCCN QA76.95 .S68 1984. US$32.50.

Cody:1984:PRW

[66] William J. Cody, Jr., Jerome T. Coonen, David M. Gay, K. Hanson, David G. Hough, William Kahan, Richard Karpinski, John F. Palmer, Frederic N. Ris, and David Stevenson. A proposed radix- and word-length-independent standard for floating-point arithmetic. IEEE Micro, 4(4):86–100, July/August 1984. CODEN IEMIDZ.

ISSN 0272-1732 (print), 1937-4143 (electronic).

Cody:1984:SEF

[67] W. J. Cody. Software for the elementary functions. Presented at the SIAM National Conference, Seattle, WA., July 1984. Cited in [?, Reference 5] in elefunt.bib and fparith.bib: was this published in print??

Cody:1984:SPS

[68] W. J. Cody. FUNPACK—A package of special function routines. In Cowell [103], pages 49–67. ISBN 0-13-823501-5. LCCN QA76.95 .S68 1984. US$32.50. Republication of [55].

Cody:1984:SSF

[69] W. J. Cody. Software for special functions. Technical Report ANS/MCS-TM-37, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, October 1984. Also published as [73].

Cody:1984:STMa

[70] W. J. Cody. Second thoughts on the mathematical software effort: a perspective. Technical Report ANL-84-83, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, October 1984. 18 pp.

Cody:1984:STMb

[71] W. J. Cody. Second thoughts on the mathematical software effort: a perspective. In Dongarra et al. [124], pages 129– 152. LCCN QA297.S879 1984. Copies of lecture slides. See [70] for the text.

Cody:1985:PRW

[72] W. J. Cody, J. T. Coonen, D. M. Gay, K. Hanson, D. Hough, W. Kahan, R. Karpinski, J. Palmer, F. N. Ris, and D. Stevenson. A proposed radixand word-length-independent standard for floating-point arithmetic. ACM SIGNUM Newsletter, 20(1):37–51, January 1985. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Cody:1985:SSF

[73] W. J. Cody. Software for special functions. In Rendiconti del Seminario Matematico, Fascicolo Speciale, Convegno su “Special Functions”, Università e Politecnico Torino, Italy, pages 91–116. ??, ??, 1985. ISBN ?? LCCN ?? Also published as [69].

Cody:1986:ALB

[74] W. J. Cody. An alternative library under 4.2 BSD UNIX on a VAX 11/780. Technical Report ANL-86-10, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, February 1986. iii + 30 pp.

Cody:1986:ETRa

[75] W. J. Cody. ELEFUNT test results under X1.4 on the Encore Multimax. Technical Report MCS-TM-68, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, April 1986. 11 pp.

Cody:1986:ETRb

[76] W. J. Cody. ELEFUNT test results under NS32000 Fortran V2.5.3 on the Sequent Balance. Technical Report MCS-TM-80, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, September 1986. 12 pp.

Cody:1986:ETRc

[77] W. J. Cody. ELEFUNT test results under FX/Fortran version 1.0 on the Alliant FX/8. Technical Report MCS-TM-78, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, ?? 1986. ?? pp.

Cody:1987:SPS

[78] W. J. Cody. SPECFUN—a portable special function package. In Wouk [104], pages 1–12. ISBN 0-89871-210-6. LCCN QA76.5 .W66 1985.

Cody:1988:AMS

[79] W. J. Cody. Algorithm 665. MACHAR: A subroutine to dynamically determine machine parameters. ACM Transactions on Mathematical Software, 14(4):303– 311, December 1988. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://www.acm.

                                        org/pubs/toc/Abstracts/0098-3500/
                                        51907.html.

Cody:1988:FSP

[80] W. J. Cody. Floating-point standards — theory and practice. In Moore [125], pages 99–107. ISBN 0-12-505630-3. LCCN QA76.9.E94 R45 1988.

Cody:1988:NVM

[81] W. J. Cody. New version of MACHAR available. ACM SIGNUM Newsletter, 23(3, 4):9–10, July–October 1988. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Cody:1988:PEPa

[82] W. J. Cody, Jr. Performance evaluation of programs related to the real gamma function. Mathematics and Computer Science Preprint MCS-P12-0988, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, September 1988. ?? pp. Published in [95].

Cody:1988:PEPb

[83] W. J. Cody, Jr. Performance evaluation of programs for the error and complementary error functions. Mathematics and Computer Science Preprint MCS-P13-0988, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, September 1988. Published in [92].

Cody:1989:AXF

[84] W. J. Cody. Algorithm XXX: Functions to support the IEEE standard for binary floating-point arithmetic. Technical Report MCS-P90-0789, Mathematics and Computer Science Division, Argonne National Laboratory, Argonne, IL, USA, July 1989. 7 + 9 (C source code) pp.

Cody:1989:ETR

[85] W. J. Cody. ELEFUNT test results using Titan Fortran under Ardent UNIX 2.0 on the Titan. Technical Report MCS-TM-129, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, March 1989. iii + 14 pp.

Cody:1989:PEP

[86] W. J. Cody and L. Stoltz. Performance evaluation of programs for certain Bessel functions. ACM Transactions on Mathematical Software, 15(1): 41–48, March 1989. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0098-3500/

62039.html. Also published as Technical Report MCS-P14-0988, Argonne National Laboratory, Argonne, IL, USA.

Cody:1989:UTS

[87] W. J. Cody and L. Stoltz. The use of Taylor series to test accuracy of function programs. Technical Report MCS-P61-0289, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, February 1989. ?? pp. Published in [96].

Cody:1990:APF

[88] W. J. Cody. Analysis of proposals for the floating-point standard. In Computer Arithmetic [126], pages 312–316. ISBN 0-8186-8945-5. LCCN QA76.9 .C62C66 1990. Reprint of [54]. The citation of that source in the book is incorrectly listed as volume 20, 1987, instead of the correct volume 14, 1981.

Cody:1990:AXS

[89] W. J. Cody. Algorithm XXX: SPEC-FUN: A portable package of special functions and test drivers. Technical Report MCS-P179-0990, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, September 1990. 10 pp.

Cody:1990:ETR

[90] W. J. Cody. ELEFUNT test results under AST Fortran V1.8.0 on the Sequent Symmetry. Technical Report MCS-TM-138, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, July 1990. ?? pp.

Cody:1990:NI

[91] W. J. Cody. The normal integral. Technical Report MCS-P189-1090, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, November 1990. 2 + 4 (C source code) pp.

Cody:1990:PEP

[92] W. J. Cody, Jr. Performance evaluation of programs for the error and complementary error functions. ACM Transactions on Mathematical Software, 16 (1):29–37, March 1990. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://www.

acm.org/pubs/citations/journals/
toms/1990-16-1/p29-cody/;   http:
//www.acm.org/pubs/toc/Abstracts/
0098-3500/77628.html.

Cody:1991:CPT

[93] W. J. Cody. CELEFUNT: A portable test package for complex elementary functions. Technical Report ANL-91/1, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, January 1991. iii + 21 pp.

Cody:1991:KAA

[94] W. J. Cody. Keynote address: Arithmetic standards: The long road. In Kornerup and Matula [127], page ix. ISBN 0-8186-9151-4 (case), 0-8186-6151-8 (microfiche), 0-7803-0187-0 (library binding). LCCN QA76.9.C62 S95 1991. URL http:

//www.acsel-lab.com/arithmetic/
arith10/papers/ARITH10_keynote.

pdf. IEEE catalog no. 91CH3015-5.

Cody:1991:PEP

[95] W. J. Cody, Jr. Performance evaluation of programs related to the real gamma function. ACM Transactions on Mathematical Software, 17(1):46– 54, March 1991. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://www.acm.

                                        org/pubs/toc/Abstracts/0098-3500/

103153.html. Preprint in [82].

Cody:1991:UTS

[96] W. J. Cody, Jr. and L. Stoltz. The use of Taylor series to test accuracy of function programs. ACM Transactions on Mathematical Software, 17(1):55–63, March 1991. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

                                        URL http://www.acm.org/pubs/toc/
                                        Abstracts/0098-3500/103154.html.

Cody:1993:ACP

[97] W. J. Cody. Algorithm 714: CELE-FUNT: A portable test package for complex elementary functions. ACM Transactions on Mathematical Software, 19 (1):1–21, March 1993. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://www.acm.

                                        org/pubs/toc/Abstracts/0098-3500/
                                        151272.html.

Cody:1993:AFS

[98] W. J. Cody and J. T. Coonen. Algorithm 722: Functions to support the IEEE standard for binary floatingpoint arithmetic. ACM Transactions on Mathematical Software, 19(4):443– 451, December 1993. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://www.acm.

                                        org/pubs/toc/Abstracts/0098-3500/
                                        168185.html.

Cody:1993:ASP

[99] W. J. Cody, Jr. Algorithm 715: SPEC-FUN: A portable FORTRAN package of special function routines and test drivers. ACM Transactions on Mathematical Software, 19(1):22–32, March 1993. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL http://www.acm.org/pubs/toc/
Abstracts/0098-3500/151273.html.

Cody:2003:AT

[100] William J. Cody. Approximation theory. In Ralston et al. [128], pages 73–76. ISBN 0-470-86412-5. LCCN QA76.15

.E48 2003. URL https://dl.acm.org/
doi/10.5555/1074100.1074131.

Fike:1967:MEP

[101] C. T. Fike. Methods of evaluating polynomial approximations in function evaluation routines. Communications of the Association for Computing Machinery, 10(3):175–178, March 1967. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark on efficiency [13].

Ralston:1976:ECS

[102] Anthony Ralston and Chester L. Meek, editors. Encyclopedia of Computer Science. Petrocelli/Charter, New York, NY, USA, 1976. ISBN 0-88405-321-0. xxviii + 1523 pp. LCCN QA76.15 .E56 1976. US$60.00.

Cowell:1984:SDM

[103] Wayne R. Cowell, editor. Sources and Development of Mathematical Software. Prentice-Hall series in computational mathematics. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1984. ISBN 0-13-823501-5. xii + 404 pp. LCCN QA76.95 .S68 1984. US$32.50.

Wouk:1987:NCE

[104] Arthur Wouk, editor. New Computing Environments: Microcomputers in Large-Scale Computing. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1987. ISBN 0-89871-210-6. x + 166 pp. LCCN QA76.5 .W66 1985.

Haigh:2004:IWJ

[105] Thomas Haigh. An interview with W. J. Cody. Computer History Museum interview., August 3–4,

2004. URL http://history.siam.
org/%5C/pdfs2/Cody_returned_SIAM.
pdf;   http://history.siam.org/
oralhistories/cody.htm;   https:/
/archive.computerhistory.org/resources/
access/text/2013/12/102746785-05-
01-acc.pdf.

Anonymous:2009:WCO

[106] Anonymous. William Cody obituary. Chicago Tribune, June 25, 2009. URL

https://www.chicagotribune.com/
obituaries/william-cody-il/#:~:
text=William%20%22Jim%22%20Cody,
%20Jr.,%20resident%20of;   https:
//www.legacy.com/obituaries/name/
william-cody-obituary?pid=128895074#
:~:text=William%20%22Jim%22%20Cody,
%20Jr.,%20resident%20of;   https:
//www.legacy.com/us/obituaries/
chicagotribune/name/william-cody-
obituary?id=2547132#:~:text=William%
20%22Jim%22%20Cody,%20Jr.,%20resident%
20of.

More:2009:JCP

[107] Jorge Moré and Cleve Moler. Jim Cody: Pioneer in mathematical software libraries. NA Digest postings, June 28,

2009. URL http://www.netlib.org/
na-digest-html/09/v09n26.html#1.

AFIPS:1967:ACP

[108] 1967 Spring Joint Computer Conference, April 18–20, Atlantic City, NJ, volume 30 of AFIPS conference proceedings. Thompson Book Co., Washington, DC, USA, 1967. LCCN TK7885.A1 J6 1967.

AFIPS:1969:ACPa

[109] 1967 Spring Joint Computer Conference, May 14–16, 1969, Boston, MA, volume 34 of AFIPS conference proceedings. AFIPS Press, Montvale, NJ, USA, 1969. LCCN TK7885.A1 J6 1969.

AFIPS:1969:ACPb

[110] 1969 Fall Joint Computer Conference, November 18–20, 1969, Las Vegas, Nevada, volume 35 of AFIPS conference proceedings. AFIPS Press, Montvale, NJ, USA, 1969. LCCN TK7885.A1 J6 1969.

Rice:1971:MS

[111] John R. Rice. Mathematical Software.

Academic Press, New York, NY, USA, 1971. ISBN 0-12-587250-X. xvii + 515 pp. LCCN QA1 .M26.

ACM:1972:PAA

[112] Proceedings of the ACM Annual Conference, August 1972, Boston. ACM Press, New York, NY 10036, USA, 1972. LCCN TK 7885 A84p 1972. Two volumes.

Hetzel:1973:PCP

[113] William C. Hetzel, editor. Program test methods: Proceedings of the Computer Program Test Methods Symposium held at the University of North Carolina, Chapel Hill, June 21–23, 1972. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1973. ISBN 0-13-729624-X. LCCN QA76.P69.

Evans:1974:SNM

[114] D. J. Evans, editor. Software for Numerical Mathematics: Proceedings of the Loughborough University of Technology Conference of the Institute of Mathematics and Its Applications held in April 1973. Academic Press, New York, NY, USA, 1974. ISBN 0-12-243750-0. LCCN QA297 .S591.

Watson:1975:NAD

[115] George A. Watson, editor. Numerical Analysis: Proceedings of the Dundee Conference on Numerical Analysis, July 1–4, 1975, volume 506 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. ISBN 3-540-07610-7 (print), 3-540-38129-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA297; QA3 .L28 no. 506.

Anonymous:1976:PIS

[116] David C. Hoaglin and Roy E. Welsch, editors. Proceedings of the Ninth Interface Symposium on Computer Science and Statistics, Harvard University, Massachusetts Institute of Technology, April 1–2, 1976. Prindle, Weber and Schmidt, Boston, MA, USA, 1976. ISBN 0-87150-237-2. LCCN QA 276 A1 I53 1976.

Cowell:1977:PMS

[117] Wayne Cowell, editor. Portability of Numerical Software Workshop, Oak Brook, Illinois, June 21–23, 1976, volume 57 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. ISBN 0-387-08446-0. LCCN QA297 .W65 1976.

IEEE:1977:PAR

[118] Proceedings, 1st Annual Rocky Mountain Symposium on Microcomputers: Systems, Software, Architecture: August 31–September 2, 1977, Fort Collins, Colorado. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1977. LCCN QA76.5 .R535 1977a.

IEEE:1980:PCS

[119] COMPCON (20th: 1980: San Francisco, CA) VLSI, New Architectural Horizons: COMPCON, Spring 80, Jack Tar Hotel, San Francisco, California, February 25–28, 1980: Digest of Papers. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1980. LCCN TK7885.A1 C53 1980.

Messina:1982:PMM

[120] P. C. Messina and A. Murli, editors.

Problems and Methodologies in Mathematical Software Production: International Seminar held at Sorrento, Italy, November 3–8, 1980, volume 142 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 0-387-11603-6 (New York), 3-540-11603-6 (Berlin). LCCN QA76.95 .P76 1982.

Mulvey:1982:EMP

[121] J. M. Mulvey, editor. Evaluating Mathematical Programming Techniques: Proceedings of a Conference Held at the National Bureau of Standards, Boulder, Colorado, January 5–6, 1981, volume 199 of Lecture Notes in Economics and Mathematical Systems. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 0-387-11495-5. LCCN QA402.5 .E94 1982.

Reid:1982:RBN

[122] J. K. Reid, editor. The Relationship Between Numerical Computation and Programming Languages: Proceedings of the IFIP TC2 Working Conference on the Relationship between Numerical Computation and Programming Languages, Boulder, Colorado, USA., 3–7 August, 1981. Elsevier North-Holland, Inc., New York, NY, USA, 1982. ISBN 0-444-86377-X. LCCN QA297 .I34 1981.

Gentle:1983:CSS

[123] James E. Gentle, editor. Computer Science and Statistics: Proceedings of the Fifteenth Symposium on the Interface, Houston, Texas, March 1983. North-Holland, Amsterdam, The Netherlands, 1983. ISBN 0-444-86688-4. LCCN QA276.4 .S95 1983.

Dongarra:1984:IPS

[124] Jack Dongarra, Gene Golub, Jorge Moré, and Danny Sorensen, editors. Informal proceedings of the Symposium on Computational Mathematics — State of the Art: held at Argonne National Laboratory, September 20–21, 1984, in honor of James H. Wilkinson. Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, December 1984. LCCN QA297.S879 1984. Technical Report MCS-TM-42. It consists of copies of lecture slides from the ten symposium talks.

Moore:1988:RCR

[125] Ramon E. Moore, editor. Reliability in Computing: the Role of Interval Methods in Scientific Computing, volume 19 of Perspectives in computing. Academic Press, New York, NY, USA, 1988. ISBN 0-12-505630-3. xv + 428 pp. LCCN QA76.9.E94 R45 1988.

Swartzlander:1990:CAb

[126] Earl E. Swartzlander, Jr. Computer Arithmetic, volume 2. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. ISBN 0-8186-8945-5. ix + 396 pp. LCCN QA76.9 .C62C66 1990. This is part of a two-volume collection of influential papers on the design of computer arithmetic. See also [?].

Kornerup:1991:PIS

[127] Peter Kornerup and David W. Matula, editors. Proceedings: 10th IEEE Symposium on Computer Arithmetic: June 26–28, 1991, Grenoble, France. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. ISBN 0-8186-9151-4 (case), 0-8186-6151-8 (microfiche), 0-7803-0187-0 (library binding). LCCN QA76.9.C62 S95 1991. IEEE catalog no. 91CH3015-5.

Ralston:2003:ECS

[128] Anthony Ralston, Edwin D. Reilly, and David Hemmendinger, editors. Encyclopedia of Computer Science. Wiley, New York, NY, USA, fourth edition, 2003. ISBN 0-470-86412-5. xxix + 2034 pp. LCCN QA76.15 .E48 2003. URL http:/

/www.e-streams.com/es0707/es0707\
_3357.htm;   http://www.loc.gov/
catdir/bios/wiley046/2003283283.

                                        htm;   http://www.loc.gov/catdir/
                                        description/wiley041/2003283283.
                                        htm;   http://www.loc.gov/catdir/
                                        toc/wiley041/2003283283.htm.